OpenStudy (anonymous):

2y=-x+5 is the equation of a line that passes thru tge point (1,2) and has the slope of -1/2. (True or False)?

OpenStudy (johnweldon1993):

turn this into slope intercept form y = mx + b turn your equation into one that looks like this *hint, notice how y is alone in the equation you want*

OpenStudy (anonymous):

So true would be right?

OpenStudy (johnweldon1993):

correct!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I have false for y=-3x+4 is an equation that represents a line parallel to the line 3x+2y=9. would false be correct?

OpenStudy (johnweldon1993):

correct!

OpenStudy (anonymous):

one more lol! I have NO as the answer. Does the graph of the straight line with slope of -5 and y- intercept of -8 pass through the point (2,-3)

OpenStudy (johnweldon1993):

sorry for the wait....again...correct!

OpenStudy (anonymous):

Also I have NO for Does the line y+4x=7 pass through the point (1,3)?

OpenStudy (johnweldon1993):

lol what happened to 1 more :P lets check this one over again (x,y) plug your x and y into that equation again

OpenStudy (anonymous):

sorry haha.

OpenStudy (johnweldon1993):

what do you have for an answer now?

OpenStudy (anonymous):

I have No.

OpenStudy (johnweldon1993):

alright lets do this one out y + 4x = 7 point (1,3) 3 + 4(1) = 7 3 + 4 = 7 7 = 7? it most certainly does....so this point IS on this line

OpenStudy (anonymous):

Well i'm wrong I guess thank you!

OpenStudy (johnweldon1993):

not a problem!